6.40 kJ of energy is needed to raise the temperature of a calorimeter and its contents 1.00 oC. When 2.00 g of propyne (C3H4(g)) is burned in a calorimeter, the temperature of the calorimeter was raised 15.0 oC. What is the molar heat of combustion of propyne?

(Please note that you must include units throughout all of your calculations and provide your answer in scientific notation to the correct number of significant digits.)
1
Expert's answer
2019-08-13T03:20:42-0400

According to the combustion equation:

Q1 = ΔHcomb × N = ΔHcomb × (mpropyne / Mrpropyne),

where Q1 - heat of propyne combustion, ΔHcomb - molar heat of combustion, N - number of moles, mpropyne - propyne mass, Mrpropyne - propyne molecular weight.


According to the heat equation:

Q2 = mcal × cp × ΔT,

where Q2 - heat absorbed by a calorimeter and its contents, mcal - mass of a calorimeter and its contents, cp - specific heat of calorimeter and its content, ΔT - temperature change.


As Q1 = Q2:

ΔHcomb × (mpropyne / Mrpropyne) = mcal × cp × ΔT.

From here:

ΔHcomb = (Mrpropyne × mcal × cp × ΔT) / mpropyne,

where Mrpropyne = 40 g/mol, ΔT = 15°C, mpropyne = 2 g.


As Q0 = mcal × cp × ΔT1°C, mcal × cp = Q0 / ΔT1°C,

where Q0 - energy needed to raise the temperature of a calorimeter and its contents 1 °C.


As a result:

ΔHcomb = [Mrpropyne × (Q0 / ΔT1°C)× ΔT] / mpropyne = [40 g/mol × 6.4 kJ/°C × 15°C] / 2 g = 1920 kJ/mol.


Answer: -1920 kJ/mol.